Two men at P and Q are in front of a cliff. P is 340 m away from the cliff and Q is 340 m away from P. The man at P sounds a whistle continuously for half a second. What is the direction of hearing the sound of echo by P? What is the time gap between the two sounds that Q hears?

Grade:9

1 Answers

Sanjeev Soneji
6 years ago
1.Towards Himself.2.solutionAssuming velocity of sound be 340 msLet the first whistle heard by Q at time T1Let the second whistle heard(echo) be heard at T2T1=340/340=1 secT2=3x340/340=3 secAlso we would also consider the half second of whistle heard Therefore gap would be 3-1-1/2=3/2 sec